The value of Kc for the reaction of phosphorus pentachloride to make phosphorus trichloride and chlorine is 2.2915. The concentration of phosphorus pentachloride 1.34427 M with no products. What is the equilibrium concentration (in M) of chlorine?
Answer #1

Pily & de Pdg - 1.34427 2 1.34427-2 K= [pl,780, [pels I 2.2915 = x2 (1.34427-2 art 2.24150 - 3.0804 = 0 xi 950 2 3 m [ca, 3 =